OpenStudy (anonymous):

I'm not sure if I'm right, I will award medals! May someone check my answer? I got B. @ganeshie8

OpenStudy (anonymous):

3rd one

ganeshie8 (ganeshie8):

your'e right ! would be nice to see your work though :)

OpenStudy (anonymous):

ganeshie8 can you help me on a problem please

OpenStudy (anonymous):

That's what confuses me I've seen this problem before and have seen one person get option A, another person get option B, and another person get option C. (Of course, each person showed a different variation in their work.) ----------------------------- For option A: V = pi r^2 * h 50.24 = 4pi h h = 50.24/4pi V = 9pi * 50.24/4pi = 9 * 50.24/4 = 113.04 ~~~~~~~~~~~~~~~~~~~~~~~~ For option B: \[50.24 \times 27 \div 8 = \] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~` For option C: (as the another user @jimo got above) I don't remember the work for this method. @ganeshie8 . Please clear up my confusion... :/

OpenStudy (anonymous):

The correct answer is B.